# Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization

### v1

@@ Abstract
There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems. This perception may partly stem from the intuition that GPs struggle with high-dimensional inputs for covariance modeling and function estimation. While these concerns seem reasonable, empirical evidence supporting this belief is lacking. In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization. Surprisingly, the performance with standard GP consistently ranks among the best, often outperforming existing BO methods specifically designed for high-dimensional optimization by a large margin. Contrary to the stereotype, we found that standard GP can serve as a capable surrogate for learning high-dimensional target functions. Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also proves robust in accommodating various structures within the target functions. Furthermore, with standard GP, achieving promising optimization performance is possible by only using maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models. We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems.
@@ 2 Standard Bayesian Optimization
Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$. Due to the GP prior, the function values at $\bf X$}, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, $\bf K$} is the kernel matrix on $\bf X$}, and each ${\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Next, a Gauss noise model is used to model the likelihood of $\bf y$}, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance. One can marginalize $\bf f$} out,
@@ 4.1 Experimental Settings
  • DNA (180) (Šehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020). Prior analysis (Šehić et al., 2022) shows that only around 43 variables are relevant to the target function.
@@ 4.2.3 Computational Efficiency
[…]
@@ A.1 Definition of Synthetic Functions
where ${\bf x}\in[-5,5]^{D}$, and we set $c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$). The optimum is at ${\bf x}^{\dagger}=[c_{1}-2.903534,\ldots,c_{D}-2.903534]$.
@@
where ${\bf x}\in[-2.048,2.048]^{D}$. We set $c_{1},...,c_{D}]$ as an evenly spaced sequence in [2,2][-2,2], where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
@@ A.2 NN-Prune Details
Given specific per-layer pruning rates ${\bf x}=[x_{1},\ldots,x_{151}]$, the pruning procedure is as follows. The weights in each layer $i$ are first ranked according to their magnitudes, and the bottom $x_{i}$ portion of the weights are removed. The pruned network is then fine tuned on Tiny ImageNet (Le and Yang, 2015). The prediction accuracy after pruning and fine tuning on a separate test dataset is used as the target function. To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from $\bf x$}. Note that the penalty is only incurred when $\alpha>\beta$. When $\alpha<\beta$, namely, the network is pruned to be sparser than the expected level, the test accuracy typically experienced a significant decline. Consequently, the optimization process automatically adjusts to increase $\beta$. We set $\lambda=8.0$ to introduce a strong penalty. In our experiment, we employed $\alpha=0.5$.
@@ A.4 Computing Environment
[…]

### v2

@@ Abstract
There has been a long-standing and widespread belief that Bayesian Optimization (BO) with standard Gaussian process (GP), referred to as standard BO, is ineffective in high-dimensional optimization problems. This perception may partly stem from the intuition that GPs struggle with high-dimensional inputs for covariance modeling and function estimation. While these concerns seem reasonable, empirical evidence supporting this belief is lacking. In this paper, we systematically investigated BO with standard GP regression across a variety of synthetic and real-world benchmark problems for high-dimensional optimization. Surprisingly, the performance with standard GP consistently ranks among the best, often outperforming existing BO methods specifically designed for high-dimensional optimization by a large margin. Contrary to the stereotype, we found that standard GP can serve as a capable surrogate for learning high-dimensional target functions. Without strong structural assumptions, BO with standard GP not only excels in high-dimensional optimization but also proves robust in accommodating various structures within the target functions. Furthermore, with standard GP, achieving promising optimization performance is possible by only using maximum likelihood estimation, eliminating the need for expensive Markov-Chain Monte Carlo (MCMC) sampling that might be required by more complex surrogate models. We thus advocate for a re-evaluation and in-depth study of the potential of standard BO in addressing high-dimensional problems. The code for our experiment is available at https://github.com/XZT008/Standard-GP-is-all-you-need-for-HDBO.
@@ 2 Standard Bayesian Optimization
Suppose we have collected a set of training inputs ${\bf X}=[{\bf x}_{1},\ldots,{\bf x}_{N}]^{\top}$ and (noisy) outputs ${\bf y}=[y_{1},\ldots,y_{N}]^{\top}$. Due to the GP prior, the function values at ${\bf X}$, denoted as ${\bf f}=[f({\bf x}_{1}),\ldots,f({\bf x}_{N})]^{\top}$, follow a multi-variate Gaussian distribution, $p({\bf f})=\mathcal{N}({\bf f}|{\bf m},{\bf K})$, where ${\bf m}=[m({\bf x}_{1}),\ldots,m({\bf x}_{N})]^{\top}$, ${\bf K}$ is the kernel matrix on ${\bf X}$, and each $[{\bf K}]_{ij}=\kappa({\bf x}_{i},{\bf x}_{j})$. Next, a Gauss noise model is used to model the likelihood of ${\bf y}$, namely, $p({\bf y}|{\bf f})=\mathcal{N}({\bf y}|{\bf f},\sigma^{2}{\bf I})$ where $\sigma^{2}$ is the noise variance. One can marginalize ${\bf f}$ out,
@@ 4.1 Experimental Settings
  • DNA (180) (vSehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020). Prior analysis (vSehić et al., 2022) shows that only around 43 variables are relevant to the target function.
@@ 4.2.3 Computational Efficiency
Mopta08(124) 0.613$\pm$0.156 -0.198$\pm$0.233 0.584$\pm$0.158 -1.176$\pm$0.080 -1.177$\pm$0.327
Rover(60) -1.209$\pm$0.194 -0.908$\pm$0.068 -0.909$\pm$0.108 -1.270$\pm$0.094 -1.687$\pm$0.251
DNA(180) 0.005$\pm$0.175 -0.459$\pm$0.142 -0.066$\pm$0.225 -1.346$\pm$0.108 -1.572$\pm$0.432
SVM(388) 0.296$\pm$0.097 -0.090$\pm$0.065 0.160$\pm$0.113 -1.175$\pm$0.116 -1.077$\pm$0.148
Ackley(150, 150) 0.720$\pm$0.092 0.899$\pm$0.058 0.809$\pm$0.053 -0.852$\pm$0.091 -0.266$\pm$0.091
Ackley150(300, 150) 0.617$\pm$0.191 0.367$\pm$0.081 0.536$\pm$0.095 -0.870$\pm$0.097 -0.478$\pm$0.204
Hartmann6 (300, 6) 1.681$\pm$0.110 -1.031$\pm$0.117 2.152$\pm$0.971 -1.433$\pm$0.249 -1.689$\pm$0.411
Rosenbrock(100, 100) 0.602$\pm$0.069 0.172$\pm$0.099 0.553$\pm$0.084 -1.007$\pm$0.121 -0.854$\pm$0.355
Rosenbrock(300, 100) 0.436$\pm$0.098 -0.256$\pm$0.071 0.265$\pm$0.115 -1.118$\pm$0.065 -1.380$\pm$0.329
Stybtang(200, 200) 0.558$\pm$0.122 0.215$\pm$0.076 0.463$\pm$0.098 -1.050$\pm$0.099 -1.238$\pm$0.436
@@ A.1 Definition of Synthetic Functions
where ${\bf x}\in[-5,5]^{D}$, and we set $[c_{1},\dots,c_{D}]$ as an evenly spaced sequence in $[0,7.5]$ ($c_{1}=0$ and $c_{D}=7.5$). The optimum is at ${\bf x}^{\dagger}=[c_{1}-2.903534,\ldots,c_{D}-2.903534]$.
@@
where ${\bf x}\in[-2.048,2.048]^{D}$. We set $[c_{1},...,c_{D}]$ as an evenly spaced sequence in $[-2,2]$, where $c_{1}=-2$ and $c_{D}=2$. The optimum is at $x^{\dagger}=[c_{1}+1,...,c_{D}+1]$.
@@ A.2 NN-Prune Details
Given specific per-layer pruning rates ${\bf x}=[x_{1},\ldots,x_{151}]$, the pruning procedure is as follows. The weights in each layer $i$ are first ranked according to their magnitudes, and the bottom $x_{i}$ portion of the weights are removed. The pruned network is then fine tuned on Tiny ImageNet (Le and Yang, 2015). The prediction accuracy after pruning and fine tuning on a separate test dataset is used as the target function. To respect the global pruning rate constraint $\alpha$, we add to the objective a soft penalty term $-\lambda\cdot\max(\alpha-\beta,0)$, where $\beta$ is the global pruning rate reflected from ${\bf x}$. Note that the penalty is only incurred when $\alpha>\beta$. When $\alpha<\beta$, namely, the network is pruned to be sparser than the expected level, the test accuracy typically experienced a significant decline. Consequently, the optimization process automatically adjusts to increase $\beta$. We set $\lambda=8.0$ to introduce a strong penalty. In our experiment, we employed $\alpha=0.5$.
@@ A.4 Computing Environment
Mopta08(124) 0.048$\pm$0.015 0.126$\pm$0.041 0.052$\pm$0.018 0.654$\pm$0.146 0.659$\pm$0.096
Rover(60) 0.493$\pm$0.126 0.449$\pm$0.108 0.453$\pm$0.121 0.813$\pm$0.168 1.266$\pm$0.352
DNA(180) 0.098$\pm$0.028 0.175$\pm$0.070 0.115$\pm$0.039 0.827$\pm$0.173 0.955$\pm$0.335
SVM(388) 0.042$\pm$0.010 0.098$\pm$0.012 0.061$\pm$0.017 0.624$\pm$0.159 0.493$\pm$0.132
Ackley(150, 150) 0.010$\pm$0.003 0.004$\pm$0.001 0.004$\pm$0.001 0.367$\pm$0.074 0.102$\pm$0.015
Ackley(300, 150) 0.010$\pm$0.003 0.018$\pm$0.004 0.010$\pm$0.002 0.360$\pm$0.109 0.177$\pm$0.075
Hartmann6(300, 6) 0.020$\pm$0.009 0.769$\pm$0.124 0.077$\pm$0.045 1.012$\pm$0.154 1.286$\pm$0.210
Rosenbrock(100, 100) 0.049$\pm$0.006 0.124$\pm$0.020 0.062$\pm$0.008 0.601$\pm$0.095 0.807$\pm$0.248
Rosenbrock(300, 100) 0.048$\pm$0.012 0.202$\pm$0.035 0.073$\pm$0.021 0.568$\pm$0.057 1.290$\pm$0.882
Stybtang(200, 200) 0.040$\pm$0.014 0.078$\pm$0.012 0.044$\pm$0.014 0.502$\pm$0.066 0.861$\pm$0.329